# TESGNN: Temporal Equivariant Scene Graph Neural Networks for Efficient and Robust Multi-View 3D Scene Understanding

### v2

@@ VI-D Ablation Study
⓪ SGFN 37.82 48.74 62.82 $\mathbf{88.08}$ 81.41 98.22
ESGNN_1 $\mathbf{42.30}$ $\mathbf{53.30}$ $\mathbf{63.21}$ 86.70 94.34 $\mathbf{98.30}$
@@
④ ESGNN_2X 37.94 50.58 59.97 85.23 $\mathbf{94.53}$ 98.01
TABLE III: Evaluation of different ESGNN architectures on scene graph generation task on 3DSSG-l20 dataset. ② is our best performer and is used for evaluation in Section VI-C.
Models ③ and ④ perform well on the train set but poorly on the validation and test sets, potentially suffering overfitting as they contain more layers. Models ② and ④ have higher edge recalls in several initial epochs, but decline in the later epochs, as shown in Fig 5(b), 5(e).
VI-E Temporal Graph Matching Evaluation
Experiment Setup: We assess our work from two key perspectives. First, we evaluate our Graph Matching Network using different backbone scene graph extractors, including the state-of-the-art SGFN [11] and our proposed ESGNN. Second, we compare the effectiveness of the SGFN and ESGNN backbones on the state-of-the-art SG-PGM [30]. To ensure the same set-up, we reimplement and evaluate SG-PGM with our proposed dataset mentioned in Section VI-A.
@@ VI-E Temporal Graph Matching Evaluation
ESGNN $\mathbf{30.14}$ $\mathbf{46.94}$ $\mathbf{57.63}$ $\mathbf{73.86}$ $\mathbf{38.54}$ $\mathbf{42.10}$
SG-PGM SGFN 18.54 33.28 $\mathbf{45.73}$ $\mathbf{62.82}$ 25.91 30.06
(point) ESGNN $\mathbf{20.91}$ $\mathbf{33.55}$ 44.41 62.01 $\mathbf{27.23}$ $\mathbf{30.85}$
@@
ESGNN $\mathbf{70.15}$ $\mathbf{84.06}$ $\mathbf{90.09}$ $\mathbf{95.25}$ $\mathbf{78.93}$ $\mathbf{81.94}$
TABLE IV: Evaluation for Graph Matching Network with 30 training epochs. Notice that k=1 is the accuracy of the node-matching and R@1 = MRR@1
Better outcome with ESGNN backbone scene graph extractor: Table IV demonstrates that utilizing our proposed ESGNN as the backbone scene graph extractor consistently outperforms SGFN across all metrics, on both graph matching networks. This underscores the advantage of ESGNN’s symmetry-preserving property, which enhances the representation of nodes and edges in 3D point clouds, making them more resilient to transformations and significantly improving graph matching performance.
Better outcome with our Graph Matching Network: Although SG-PGM [30] achieved competitive results for node matching in their setting, it performs poorly in our scenario, where node and edge features are extracted from a scene graph extractor model rather than one-hot encoded from the ground truth. In our case, both nodes and edges have more labels, leading to several ambiguous predictions, including overlaps, incorrect assignments, and permutations. This explains why SG-PGM fails in this setting. In contrast, our model leverages the equivariant properties of nodes and edges while applying sum pooling, ensuring permutation invariance and significantly outperforming SG-PGM in this scenario, indicating the robustness of the model.
VII Conclusion

### v3

@@ 6.4 Ablation Study
⓪ SGFN 37.82 48.74 62.82 88.08 81.41 98.22
ESGNN_1 42.30 53.30 63.21 86.70 94.34 98.30
@@
④ ESGNN_2X 37.94 50.58 59.97 85.23 94.53 98.01
Table 4: Evaluation of different ESGNN architectures on the scene graph generation task using the 3DSSG-l20 dataset. ② is our best performer and is used for evaluation in Section 6.3.
Models ③ and ④ perform well on the train set but poorly on the validation and test sets, potentially suffering from overfitting as they contain more layers. Models ② and ④ have higher edge recalls in several initial epochs, but decline in the later epochs, as shown in Fig. 5(b), 5(e). To prevent overfitting, we notice that using a single EGCL layer is enough. We also experiment with different dropout rates between 0.3 and 0.5. However, a dropout rate of 0.5 led to underfitting, preventing the model from converging during training. We select dropout rate of 0.3 due to its balanced outcome.
6.5 Temporal Graph Matching Evaluation
Experiment Setup: We assess our work from two key perspectives. First, we evaluate our Graph Matching Network using different backbone scene graph extractors, including the state-of-the-art SGFN Wu et al. (2021b) and our proposed ESGNN. Second, we compare the effectiveness of the SGFN and ESGNN backbones on the state-of-the-art SG-PGM Xie et al. (2024). To ensure the same set-up, we reimplement and evaluate SG-PGM with our proposed dataset mentioned in Section 6.1. For top-K retrieval, we typically set K $\leq$ 5 to focus on the most relevant nodes. Using a larger K typically tends to inflate the score, but adds little insight in evaluating the similarity matching.
@@ 6.5 Temporal Graph Matching Evaluation
ESGNN 30.14 46.94 57.63 73.86 38.54 42.10
SG-PGM (point) SGFN 18.54 33.28 45.73 62.82 25.91 30.06
ESGNN 20.91 33.55 44.41 62.01 27.23 30.85
@@
ESGNN 70.15 84.06 90.09 95.25 78.93 81.94
Table 5: Evaluation of the Graph Matching Network with 30 training epochs. Note that R@1 = MRR@1 reflects node-matching accuracy.
Better outcome with ESGNN backbone scene graph extractor: Table 5 demonstrates that utilizing our proposed ESGNN as the backbone scene graph extractor consistently outperforms SGFN across all metrics, on both graph matching networks. This underscores the advantage of ESGNN’s symmetry-preserving property, which enhances the representation of nodes and edges in 3D point clouds, making them more resilient to transformations and significantly improving graph matching performance.
Better outcome with our Graph Matching Network: Although SG-PGM Xie et al. (2024) achieved competitive results for node matching in their setting, it performs poorly in our scenario, where node and edge features are extracted from a scene graph extractor model rather than one-hot encoded from the ground truth. In our case, both nodes and edges have more labels, leading to several ambiguous predictions, including overlaps, incorrect assignments, and permutations. This explains why SG-PGM fails in this setting. In contrast, our model leverages the equivariant properties of nodes and edges while applying sum pooling, ensuring permutation invariance and significantly outperforming SG-PGM in this scenario, indicating the robustness of the model.
7 Conclusion
@@
Appendix A GitHub
We provide the GitHub Repository  [^3] Link: https://github.com/HySonLab/TESGraph/tree/main for our TESGNN implementation.
Appendix B Equivariance of ESGNN
This section demonstrates that our model is translation equivariant with respect to $\mathbf{x}$ for any translation vector $g\in\mathbb{R}^{n}$, and rotation and reflection equivariant with respect to $\mathbf{x}$ for any orthogonal matrix $Q\in\mathbb{R}^{n\times n}$. Formally, we prove that the model satisfies:
$Q\mathbf{x}^{l+1}+g,\mathbf{h}^{l+1}=\operatorname{ESGNN}\left(Q\mathbf{x}^{l}+g,\mathbf{h}^{l}\right)$
B.1 FAN-GCL Layer
Node Update:
$\mathbf{h}_{i}^{\ell+1}=g_{v}\left(\left[\mathbf{h}_{i}^{\ell},\max_{j\in\mathcal{N}(i)}\left(\operatorname{FAN}\left(\mathbf{h}_{i}^{\ell},\mathbf{e}_{ij}^{\ell},\mathbf{h}_{j}^{\ell}\right)\right)\right]\right),$
Edge Update:
$\mathbf{e}_{ij}^{\ell+1}=g_{e}\left(\left[\mathbf{h}_{i}^{\ell},\mathbf{e}_{ij}^{\ell},\mathbf{h}_{j}^{\ell}\right]\right),$
Equivariance:
Assuming $\mathbf{h}^{\ell}$ is invariant to $\mathrm{E}(n)$ transformations on $\mathbf{x}$—i.e., no information about the absolute position or orientation of $\mathbf{x}^{\ell}$ is encoded into $\mathbf{h}^{\ell}$—then the outputs $\mathbf{h}_{i}^{\ell+1}$ and $\mathbf{e}_{ij}^{\ell+1}$ of FAN-GCL are also invariant.
B.2 EGCL Layer
Node Update:
$h_{i}^{(l+1)}=h_{i}^{(l)}+g_{v}\left(\text{concat}\left(h_{i}^{(l)},\sum_{j\in\mathcal{N}(i)}e_{ij}^{(l)}\right)\right),$
Edge Update:
$e_{ij}^{(l+1)}=g_{e}\left(\text{concat}\left(h_{i}^{(l)},h_{j}^{(l)},\|\mathbf{x}_{i}^{(l)}-\mathbf{x}_{j}^{(l)}\|^{2},e_{ij}^{(l)}\right)\right),$
Coordinate Update:
$\mathbf{x}_{i}^{(l+1)}=\mathbf{x}_{i}^{(l)}+\sum_{j\in\mathcal{N}(i)}(\mathbf{x}_{i}^{(l)}-\mathbf{x}_{j}^{(l)})\cdot\phi_{\text{coord}}(e_{ij}^{(l)}),$
Equivariance:
Since $\mathbf{h}^{\ell}$ is invariant to $\mathrm{E}(n)$ transformations of $\mathbf{x}$, the output $e_{ij}^{(l+1)}$ of EGCL is also invariant because the distance between two particles is invariant to translations:
$\left\|\mathbf{x}_{i}^{l}+g-\left[\mathbf{x}_{j}^{l}+g\right]\right\|^{2}=\left\|\mathbf{x}_{i}^{l}-\mathbf{x}_{j}^{l}\right\|^{2},$
and invariant to rotations and reflections:
$\left\|Q\mathbf{x}_{i}^{l}-Q\mathbf{x}_{j}^{l}\right\|^{2}$ $=\left(\mathbf{x}_{i}^{l}-\mathbf{x}_{j}^{l}\right)^{\top}Q^{\top}Q\left(\mathbf{x}_{i}^{l}-\mathbf{x}_{j}^{l}\right)$
$=\left(\mathbf{x}_{i}^{l}-\mathbf{x}_{j}^{l}\right)^{\top}\mathbf{I}\left(\mathbf{x}_{i}^{l}-\mathbf{x}_{j}^{l}\right)$
$=\left\|\mathbf{x}_{i}^{l}-\mathbf{x}_{j}^{l}\right\|^{2}$
Thus, the edge operation becomes invariant:
$e_{ij}^{(l+1)}$ $=g_{e}\left(\mathbf{h}_{i}^{l},\mathbf{h}_{j}^{l},\left\|Q\mathbf{x}_{i}^{l}+g-\left[Q\mathbf{x}_{j}^{l}+g\right]\right\|^{2},e_{ij}^{(l)}\right)$
$=g_{e}\left(\mathbf{h}_{i}^{l},\mathbf{h}_{j}^{l},\left\|\mathbf{x}_{i}^{l}-\mathbf{x}_{j}^{l}\right\|^{2},e_{ij}^{(l)}\right)$
The coordinates $\mathbf{x}$ are also $\mathrm{E}(n)$ equivariant. We prove this by showing that an $\mathrm{E}(n)$ transformation of the input yields the same transformation of the output. Since $e_{ij}^{(l+1)}$ is already invariant, we aim to show:
$Q\mathbf{x}_{i}^{l+1}+g$
$=Q\mathbf{x}_{i}^{l}+g+C\sum_{j\neq i}\left(Q\mathbf{x}_{i}^{l}+g-\left[Q\mathbf{x}_{j}^{l}+g\right]\right)g_{coord}\left(e_{ij}^{(l+1)}\right)$
We have the following derivation:
$Q\mathbf{x}_{i}^{l}+g+C\sum_{j\neq i}\left(Q\mathbf{x}_{i}^{l}+g-Q\mathbf{x}_{j}^{l}-g\right)g_{coord}\left(e_{ij}^{(l+1)}\right)$
$=Q\mathbf{x}_{i}^{l}+g+QC\sum_{j\neq i}\left(\mathbf{x}_{i}^{l}-\mathbf{x}_{j}^{l}\right)g_{coord}\left(e_{ij}^{(l+1)}\right)$
$=Q\left(\mathbf{x}_{i}^{l}+C\sum_{j\neq i}\left(\mathbf{x}_{i}^{l}-\mathbf{x}_{j}^{l}\right)g_{coord}\left(e_{ij}^{(l+1)}\right)\right)+g$
$=Q\mathbf{x}_{i}^{l+1}+g$
Thus, a transformation $Q\mathbf{x}^{l}+g$ on $\mathbf{x}^{l}$ leads to the same transformation on $\mathbf{x}^{l+1}$, while $\mathbf{h}^{l+1}$ remains invariant, satisfying:
$Q\mathbf{x}^{l+1}+g,\mathbf{h}^{l+1}=\operatorname{EGCL}\left(Q\mathbf{x}^{l}+g,\mathbf{h}^{l}\right).$
Both FAN-GCL and EGCL layers are designed to maintain equivariance under translation, rotation, and permutation transformations, making them well-suited for tasks requiring spatial and relational reasoning in graph-based neural networks.
Appendix C Towards Temporal Model’s Training and Future Works
C.1 Loss Selection and Training With Multiple Losses
This section explores various training strategies and loss functions applicable to our Temporal Similarity Matching. Our Temporal Model adopts the training objectives of embedding models, which are to optimize the embedding model such that distances between similar embeddings are minimized, while those between dissimilar embeddings are maximized. There are several common training strategies  [^4] Sentence Transformer Docs: We mostly refer to this well-scripted documentation for loss selection as well as different training strategies. Although its main application is for text embedding, it is very helpful to our proposed approach. Reimers & Gurevych (2019) as mentioned below.
Contrastive Loss:
This loss expects an embedding pair and a label of either 0 or 1. For label 1, the distance between the two embeddings is reduced. Otherwise, the distance between the embeddings is increased.
Online Contrastive Loss:
This leverages the Contrastive Loss, but computes the loss only for pairs of hard positive (positives that are far apart) or hard negative (negatives that are close) per iteration. This loss empirically produces better performances. In our training setting for the Temporal Model, we utilize this as our main loss.
Multiple Negatives Ranking Loss:
This loss expects a batch consisting of embedding pairs $(a_{1},p_{1}),(a_{2},p_{2})...,(a_{n},p_{n})$ where we assume that $(a_{i},p_{i})$ are a positive pair and $(a_{i},p_{j}),i\neq j$ are negative pairs. It then minimizes the negative log-likelihood for softmax normalized scores. To improve the robustness of the model, we can also provide multiple hard negative pairs per batch during training.
Cosine Similarity Loss:
A fundamental approach. In this loss, the similarity of embeddings is computed with Cosine Similarity and compared (by Mean-Squared Error) to the labeled similarity score. In our case, we can treat label 0 (dissimilar pairs) and label 1 (similar pairs) like a float similarity score. However, this loss often takes a significantly longer time to compute due to the cosine similarity calculation.
Training strategy with multiple losses:
In this strategy, we calculate each loss and update the model weights iteratively per training iteration. An example setup is to calculate the Contrastive Loss followed by the Multiple Negatives Ranking Loss. This strategy often enhances the outcome model and provides better results. In future work, we will improve our Temporal Model by applying this strategy.
C.2 Extended Discussion on Point Cloud Data Preprocessing
As our approach is built upon segmentation of 3D point cloud data, we assume that the segmentation results are reasonably accurate. Nonetheless, it is important to emphasize that the preprocessing stage in segmentation should not be overlooked. Recent studies, such as Nguyen et al. (2024b), propose lightweight methods to refine segmentation inaccuracies through non-parametric statistical approaches. Meanwhile, Grcić et al. (2023) employs CNN-based architectures that encode RGB and depth images separately and utilize their extracted features to correct errors in RGB-based segmentation. We suggest that such denoising strategies can substantially improve segmentation quality, thereby positively influencing the overall performance of scene graph generation.
Appendix D Implementation Details for ESGNN
D.1 Network Architecture
The setup for the ESGNN architecture is provided in the code block below. Our output node and edge dimensions are set to be 256. As discussed in Section 6.3. Ablation Study of the main paper, we experimented with different layer settings for the training. Each ESGNN layer is defined as a combination of 1 FAN-GCL layer followed by 1 EGCL layer to produce the final node and edge embeddings. For the FAN-GCL, we use 8 Attention heads.
model:
node_feature_dim: 256
edge_feature_dim: 256
gnn:
hidden_dim: 256
num_layers: 1
num_heads: 8
drop_out: 0.3
D.2 Training Details
We adopt a similar training strategy to state-of-the-art works 3DSSG Wald et al. (2020) and Scene Graph Fusion Wu et al. (2021b). This ensures the results of our method ESGNN, are comparable to these existing results. Our params setup is as follows:
training:
max_epoch: 200
lr: 1e-4
patient: 30
optimizer: adamw
amsgrad: true
lambda_mode: dynamic # calculate the ratio of the number of node and edge.
lambda_node: 0.1 # learning rate ratio
lambda_edge: 1.0 # learning rate ratio
scheduler:
method: reduceluronplateau
args: {gamma: 0.5,
factor: 0.9}
model_selection_metric: iou_node_cls
# Select and save best model based on the IoU
model_selection_mode: maximize
# can be maximize or minimize.
# e.g. minimize if "loss", maximize if "accuracy" or "IOU"
metric_smoothing:
method: ema
eval:
topK: 10 # Evaluate with topk = 1, 3, 5, 10, etc.
Figure 9: Training configuration for ESGNN.
Following Wu et al. (2021b), we leverage the AdamW optimizer with Amsgrad and an adaptive learning rate inversely proportional to the logarithm of the number of edges. Given a training batch containing $n$ edges and a base learning rate $\operatorname{lr}_{\text{base}}=1\times 10^{-3}$, the learning rate is adjusted as follows:
$\operatorname{lr}=\operatorname{lr}_{\text{base}}\cdot\frac{1}{\ln n}.$
We set the maximum training epoch to 200, with a learning rate of 0.0001.
[…]